# On the Localization of the Christoffel-Darboux Kernel

### v1

@@
Abstract. We establish that the Christoffel-Darboux (or Bergman) kernel localizes along the diagonal for non-pluripolar measures. This leads to several applications in the theory of Toeplitz operators, particularly concerning their asymptotic algebra property and spectral equidistribution.
@@ 1 Introduction
The primary goal of this paper is to establish a diagonal localization phenomenon for the Christoffel-Darboux kernel associated with a broad class of measures.
To state our results precisely, we fix a compact complex manifold $X$, $\dim_{\mathbb{C}}X=n$, and a big line bundle $L$ over $X$, i.e. such that for some $c>0$, we have $\dim H^{0}(X,L^{\otimes k})\geq ck^{n}$, for $k\in\mathbb{N}$ big enough. We fix a continuous Hermitian metric $h^{L}$ on $L$ and a non-pluripolar Borel measure $\mu$ on $X$, i.e. such that for any pluripolar subset $E\subset X$, we have $\mu(E)=0$. Recall that a subset $E\subset X$ is called pluripolar if in each local chart we have $E\subset\{\phi=-\infty\}$ for some plurisubharmonic (psh) function $\phi$. We denote by ${\textrm{Hilb}}_{k}(h^{L},\mu)$ the positive semi-definite form on $H^{0}(X,L^{\otimes k})$ defined for arbitrary $s_{1},s_{2}\in H^{0}(X,L^{\otimes k})$ as follows
@@
Remark that since $\mu$ is non-pluripolar, the above form is positive definite.
We assume that $k$ is large enough so that $n_{k}>0$, where $n_{k}\mathrel{\mathop{\ordinarycolon}}=\dim H^{0}(X,L^{\otimes k})$. The central object of this article is the Christoffel-Darboux (or Bergman) kernel, [1], defined as
@@
To grasp the intuitive meaning behind the Christoffel-Darboux kernel, we denote by $\Sigma_{k}$ the base loci of $L^{k}$, i.e. $\Sigma_{k}\mathrel{\mathop{\ordinarycolon}}=\{x\in X\mathrel{\mathop{\ordinarycolon}}s(x)=0\text{ for any }s\in H^{0}(X,L^{\otimes k})\}$. For $x\in X\setminus\Sigma_{k}$, we denote by $s_{k,x}\in H^{0}(X,L^{\otimes k})$ the peak section at $x$ with respect to the scalar product ${\textrm{Hilb}}_{k}(h^{L},\mu)$; recall that this means that $s_{k,x}$ is the unique section (up to a unimodular complex multiplication) which maximizes the ratio $|s(x)|_{(h^{L})^{k}}/\|s\|_{{\textrm{Hilb}}_{k}(h^{L},\mu)}$ among all $s\in H^{0}(X,L^{\otimes k})\setminus\{0\}$.
Directly from the definition of the Christoffel-Darboux kernel and the fact that it doesn’t depend on the choice of an orthonormal basis, we deduce
@@
For certain “regular” measures, this heuristic holds in a strong form; see Remark 1.2 for a brief survey. In general, however, naive versions of the phenomenon break down: one does not have pointwise convergence to zero [14, §2], nor even weak convergence of the norm at individual points in the support of $\mu$ [9]; see also Section 6. The main result of this article shows, nevertheless, that the heuristic can be recovered in an averaged form, expressed via the Christoffel-Darboux kernel.
@@
$B_{k}^{{\rm{ort}}}(x,y)\mathrel{\mathop{\ordinarycolon}}=\sum_{i=0}^{k-1}p_{i}(x)\cdot\overline{p_{i}(y)},\qquad x,y\in\mathbb{C}.$ (1.4)
To realize this Christoffel-Darboux kernel as a special case of the preceding construction, we consider the projective space $X\mathrel{\mathop{\ordinarycolon}}=\mathbb{P}^{1}$ and the hyperplane bundle $L=\mathscr{O}(1)$. We view $\mathbb{C}$ as an affine chart in $\mathbb{P}^{1}=\mathbb{C}\cup\{\infty\}$ with the standard holomorphic coordinate $z$, and consider a non-zero section $\sigma\in H^{0}(X,L)$ vanishing at $\infty$. The division by $\sigma^{k}$ gives an isomorphism between $H^{0}(X,L^{\otimes k})$ and the space of polynomials of degree $\leq k-1$ in $z$. If we now endow $L$ with a continuous metric $h^{L}$, such that $|\sigma(x)|_{h^{L}}=1$ for all $x\in K$, then the $L^{2}$-norm (1.1) coincides, under the above identification, with the $L^{2}(\mu)$-norm on the space of polynomials. In particular, (1.2) then specializes to (1.4).
@@
It follows immediately from (1.2) that each $\mu_{k}^{\rm{Berg}}$ is a probability measure. The following result confirms the heuristics above, and it is the main result of this paper.
@@
For any non-pluripolar Borel measure $\mu$, the corresponding measures $\mu_{k}^{\rm{Berg}}$ do not asymptotically place mass away from the diagonal. In other words, for any compact subset $K\subset X\times X$, not intersecting the diagonal, we have
$\lim_{k\to\infty}\iint_{K}\mu_{k}^{\rm{Berg}}=0.$ (1.6)
@@
There is a large body of work which establishes versions of Theorem 1.1 for some specific measures. A comprehensive survey would take us too far afield, so we merely cite a few representative results. In the complex geometric setting – where the measure arises from a volume form – see the works of Dai-Liu-Ma [12], Christ [10], Ma-Marinescu [25], Berman [2]. For measures supported on totally real algebraic submanifolds, see Berman-Ortega-Cerdá [4]. In the context of orthogonal polynomials, we refer to Nevai [26], Lubinsky [22], Breuer-Last-Simon [9].
@@
We note that while Theorem 1.1 appears to be new even in the context of orthogonal polynomials, the proof of it simplifies considerably in this special case (only Section 5 is relevant), and in fact a stronger statement holds under weaker assumptions: according to Remark 5.2, for arbitrary compactly supported measures $\mu$ on $\mathbb{C}$ with non-finite support and a compact $K\subset\mathbb{C}\times\mathbb{C}$, not intersecting the diagonal, there is $C>0$, so that in the notations of (1.4), for any $k\in\mathbb{N}$, we have
$\iint_{K}\big{|}B_{k}^{{\rm{ort}}}(x,y)\big{|}^{2}\cdot d\mu(x)d\mu(y)\leq C.$ (1.7)
For Bernstein-Markov measures (see [7] for a survey on the Bernstein-Markov coondition), Theorem 1.1 was recently established by the author in [14], building on earlier work of Berman-Boucksom-Witt Nyström [3], cf. also Bloom-Levenberg [5], [6], concerning the equidistribution of the diagonal Christoffel-Darboux measures (1.9), and thereby resolving a question left open by Zelditch [30] concerning the weak convergence of $\mu_{k}^{\rm{Berg}}$. The main objective of this article is to demonstrate that, by employing alternative techniques, it is possible to bypass the equidistribution result [3] and thus work within strictly weaker assumptions.
@@
Our primary motivation for Theorem 1.1 lies in the theory of Toeplitz operators. Let us recall the necessary notations: for $f\in\mathscr{C}^{0}(X)$, $k\in\mathbb{N}^{*}$, we define the Toeplitz operator $T_{k}(f)\in{\text{End}{(H^{0}(X,L^{\otimes k}))}}$ as $T_{k}(f)\mathrel{\mathop{\ordinarycolon}}=B_{k}\circ M_{k}(f)$, where $B_{k}\mathrel{\mathop{\ordinarycolon}}\mathscr{C}^{0}(X,L^{\otimes k})\to H^{0}(X,L^{\otimes k})$ is the orthogonal (Bergman) projection to $H^{0}(X,L^{\otimes k})$ with respect to (1.1), and $M_{k}(f)\mathrel{\mathop{\ordinarycolon}}H^{0}(X,L^{\otimes k})\to\mathscr{C}^{0}(X,L^{\otimes k})$ is the multiplication map by $f$.
@@
One of the key features of Toeplitz operators is their asymptotic algebra property: for $f,g\in\mathscr{C}^{0}(X)$, the composition $T_{k}(f)\circ T_{k}(g)$ is asymptotically close (with respect to some norm) to $T_{k}(f\cdot g)$, as $k\to\infty$. This phenomenon appears to have been first observed by Grenander-Szegő in [17, §7.4] in the context of certain orthogonal polynomials. It was later established in the complex-geometric setting by Bordemann-Meinrenken-Schlichenmaier [8]; see also Ma-Marinescu [24]. In the more general setting of measures verifying the Bernstein-Markov assumption, the result was recently established by the author in [14, §3]. As we shall see below, and somewhat unexpectedly, even the Bernstein-Markov assumption can in fact be dispensed with.
To describe this precisely, it is known that a version of the above asymptotic algebra property for Toeplitz operators is a consequence of the localization of the Christoffel-Darboux kernel, cf. [14, §3]. Repeating the proof of [14, Theorem 1.5], with [14, Theorem 2.1] replaced by the more general statement, Theorem 1.1, immediately yields the following result.
Corollary 1.3.
@@
For any non-pluripolar measure $\mu$, the space of Toeplitz operators verifies the asymptotic algebra property. In other words, for any $f,g\in\mathscr{C}^{0}(X)$, $p\in[1,+\infty[$, we have
@@
a) It would be interesting to classify those measures for which the $p$-Schatten norm above could be replaced by the operator norm.
b) Combining the above with (1.7), we find that in the setting of orthogonal polynomials, (1.8) holds under weaker assumptions: for arbitrary measures with non-finite support.
As an application of Corollary 1.3, we show in Section 7 that the spectral measures of Toeplitz operators equidistribute if and only if the associated diagonal Christoffel-Darboux measures (1.9) do. This result refines several classical versions of Szegő’s first limit theorem [28], as Boutet de Monvel-Guillemin [20]; see also Ma-Marinescu [23].
To describe the organization of this article, we briefly survey the main steps of the proof of Theorem 1.1. As in the earlier work [14], the key idea – albeit perhaps counterintuitive – is to investigate the off-diagonal behavior of Christoffel-Darboux kernel by analyzing its diagonal behaviour. For this, we define the diagonal Christoffel-Darboux measures on $X$ as
$\mu_{k}^{\rm{Berg},\Delta}\mathrel{\mathop{\ordinarycolon}}=\frac{1}{n_{k}}|B_{k}(x,x)|\cdot d\mu(x).$ (1.9)
Beyond this, our approach departs significantly from [14]: in the absence of the Bernstein-Markov property, little can be said about the equidistribution of (1.9), a feature that was crucial in loc. cit.
To remedy this, we establish in Sections 2 and 3, that any subsequential limit of the sequence of measures $\mu_{k}^{\rm{Berg},\Delta}$, $k\in\mathbb{N}$, assigns zero mass to subvarieties. Using this, by relying on the estimates of Lubinsky [22], we then show in Section 4 that it suffices to establish Theorem 1.1 for measures with support outside of a fixed subvariety. The proof will then be completed in Section 5, where we establish that for measures supported away from some divisors on $X$, a more precise version of Theorem 1.1 holds.
In Section 6, we discuss a pointwise version of our localization result in connection with [9, Conjecture 1.4]. In Section 7, we discuss the applications to the theory of Toeplitz operators.
Acknowledgement. The author thanks Norm Levenberg for several related discussions and Robert Berman for drawing attention to [4]. This work was supported by the CNRS, École Polytechnique, and in part by the ANR projects QCM (ANR-23-CE40-0021-01), AdAnAr (ANR-24-CE40-6184) and STENTOR (ANR-24-CE40-5905-01).
2 Christoffel-Darboux measures put no mass on subvarieties
We retain the notation from Introduction and assume, from now on and throughout the rest of the article, that the measure $\mu$ is non-pluripolar.
The primary objective of this section is to demonstrate that every subsequential limit of the diagonal Christoffel-Darboux measures vanishes on subvarieties.

### v2

# MnLargeSymbols’164 MnLargeSymbols’171

@@
Abstract. For a compact complex manifold endowed with a big line bundle and a Radon measure, we study the localization phenomena of the associated Bergman (or Christoffel-Darboux) kernel. For Bernstein-Markov measures, this results in the determination of the limiting off-diagonal Bergman measure, thereby confirming a conjecture of Zelditch. We then turn to applications in the theory of Toeplitz operators, showing in particular that they form an algebra under composition. Building on this, we then show that for Bernstein-Markov measures, the spectrum of Toeplitz operators equidistributes.
@@ 1 Introduction
The goal of this paper is to study a localization phenomenon for the Bergman kernel and to apply this study to the theory of Toeplitz operators.
To state our results precisely, we fix a compact complex manifold $X$, $\dim_{\mathbb{C}}X=n$, and a big line bundle $L$ over $X$, i.e. such that for some $c>0$, we have $\dim H^{0}(X,L^{\otimes k})\geq ck^{n}$, for $k\in\mathbb{N}$ big enough. Recall that a subset $E\subset X$ is called pluripolar if in each local chart we have $E\subset\{\phi=-\infty\}$ for some plurisubharmonic (psh) function $\phi$. We fix a continuous Hermitian metric $h^{L}$ on $L$ and a Borel measure $\mu$ on $X$ which does not give full mass to pluripolar subsets, i.e. such that for any pluripolar subset $E\subset X$, we have $\mu(E)\neq\mu(X)$. This condition is weaker than requiring $\mu$ to be non-pluripolar, i.e. such that for any pluripolar subset $E\subset X$ we have $\mu(E)=0$. We denote by ${\textrm{Hilb}}_{k}(h^{L},\mu)$ the positive semi-definite form on $H^{0}(X,L^{\otimes k})$ defined for arbitrary $s_{1},s_{2}\in H^{0}(X,L^{\otimes k})$ as follows
@@
Remark that since $\mu$ does not give full mass to analytic subsets (as every analytic subset is pluripolar), the above form (1.1) is positive definite.
We assume that $k$ is large enough so that $n_{k}>0$, where $n_{k}\mathrel{\mathop{\ordinarycolon}}=\dim H^{0}(X,L^{\otimes k})$. The central object of this article is the Bergman (or Christoffel-Darboux) kernel, [4], defined as
@@
To grasp the intuitive meaning behind the Bergman kernel, we denote by $\Sigma_{k}$ the base loci of $L^{k}$, i.e. $\Sigma_{k}\mathrel{\mathop{\ordinarycolon}}=\{x\in X\mathrel{\mathop{\ordinarycolon}}s(x)=0\text{ for any }s\in H^{0}(X,L^{\otimes k})\}$. For $x\in X\setminus\Sigma_{k}$, we denote by $s_{k,x}\in H^{0}(X,L^{\otimes k})$ the peak section at $x$ with respect to the scalar product ${\textrm{Hilb}}_{k}(h^{L},\mu)$; recall that this means that $s_{k,x}$ is the unique section (up to a unimodular complex multiplication) which maximizes the ratio $|s(x)|_{(h^{L})^{k}}/\|s\|_{{\textrm{Hilb}}_{k}(h^{L},\mu)}$ among all $s\in H^{0}(X,L^{\otimes k})\setminus\{0\}$.
Directly from the definition of the Bergman kernel and the fact that it doesn’t depend on the choice of an orthonormal basis, we deduce
@@
For certain “regular” measures, the above heuristic holds in a strong form; see Remark 1.2 for a brief survey. In general, however, naive versions of the phenomenon break down: one does not have pointwise convergence to zero, see Section 2.2, nor even weak convergence of the norm at individual points in the support of $\mu$, see Breuer-Last-Simon [19, §3-5]. The main result of this article shows, nevertheless, that the heuristic can be recovered in an averaged form.
@@
$B_{k}^{{\rm{ort}}}(x,y)\mathrel{\mathop{\ordinarycolon}}=\sum_{i=0}^{k}p_{i}(x)\cdot\overline{p_{i}(y)},\qquad x,y\in\mathbb{C}.$ (1.4)
To realize this Christoffel-Darboux kernel as a special case of the Bergman kernel, we consider the projective space $X\mathrel{\mathop{\ordinarycolon}}=\mathbb{P}^{1}$ and the hyperplane bundle $L=\mathscr{O}(1)$. We view $\mathbb{C}$ as an affine chart in $\mathbb{P}^{1}=\mathbb{C}\cup\{\infty\}$ with the standard holomorphic coordinate $z$, and consider a non-zero section $\sigma\in H^{0}(X,L)$ vanishing at $\infty$. The division by $\sigma^{k}$ gives an isomorphism between $H^{0}(X,L^{\otimes k})$ and the space of polynomials of degree $\leq k$ in $z$. If we now endow $L$ with a continuous metric $h^{L}$, such that $|\sigma(x)|_{h^{L}}=1$ for all $x\in K$, then the $L^{2}$-norm (1.1) coincides, under the above identification, with the $L^{2}(\mu)$-norm on the space of polynomials. In particular, (1.2) then specializes to (1.4).
@@
It follows immediately from (1.2) that each $\mu_{k}^{\rm{Berg}}$ is a probability measure. The following result confirms the heuristics explained after (1.3), and it is the main result of this paper.
@@
For any Borel measure $\mu$ which does not give full mass to pluripolar subsets, the corresponding measures $\mu_{k}^{\rm{Berg}}$ do not asymptotically place mass away from the diagonal. In other words, for any compact subset $K\subset X\times X$, not intersecting the diagonal, we have
$\lim_{k\to\infty}\int_{K}\mu_{k}^{\rm{Berg}}=0.$ (1.6)
@@
There is a large body of work which establishes versions of Theorem 1.1 for some specific measures. A comprehensive survey would take us too far afield, so we merely cite a few representative results. In the complex geometric setting – when the measure arises from a volume form – see the works of Dai-Liu-Ma [26], Christ [21], Ma-Marinescu [46], Berman [10]. For measures supported on totally real algebraic submanifolds, see Berman-Ortega-Cerdá [8]. In the context of orthogonal polynomials, we refer to Nevai [50], Lubinsky [41], Breuer-Last-Simon [19].
@@
We note that while Theorem 1.1 appears to be new even in the context of orthogonal polynomials, the proof of it simplifies considerably in this special case (only Section 2.4 is relevant), and in fact a stronger statement holds under weaker assumptions: according to Remark 2.9, for arbitrary compactly supported measures $\mu$ on $\mathbb{C}$ with non-finite support and a compact $K\subset\mathbb{C}\times\mathbb{C}$, not intersecting the diagonal, there is $C>0$, so that in the notations of (1.4), for any $k\in\mathbb{N}$, we have
$\iint_{K}\big|B_{k}^{{\rm{ort}}}(x,y)\big|^{2}\cdot d\mu(x)d\mu(y)\leq C,$ (1.7)
which of course refines (1.6), as $n_{k}=k+1$.
In the setting of the Bernstein-Markov measures, the above theorem can be made more precise. To explain this, let us recall some notations first. We say that $\mu$ is Bernstein-Markov with respect to $h^{L}$, if for each $\epsilon>0$, there is $C>0$, such that for any $k\in\mathbb{N}$, we have
${\textrm{Ban}}_{k}^{\infty}(K,h^{L})\leq C\cdot\exp(\epsilon k)\cdot{\textrm{Hilb}}_{k}(h^{L},\mu),\quad\text{where}\quad K={\textrm{supp}}\mu,$ (1.8)
and ${\textrm{Ban}}_{k}^{\infty}(K,h^{L})$ stands for the $L^{\infty}(K)$-norm on $H^{0}(X,L^{\otimes k})$ induced by $h^{L}$. For a survey on the Bernstein-Markov property, we recommend [11]. Here, we only mention that the Lebesgue measure on a smoothly bounded domain in $X$, or on a totally real compact submanifold of real dimension $n$, is Bernstein-Markov, see [5, Corollary 1.8], [11, Proposition 3.6] and [48, Theorem 1.3], and also the recent survey [49].
Now, we fix a non-pluripolar subset $E\subset X$. Following Siciak [56], Guedj-Zeriahi [38], we define the psh envelope $h^{L}_{E}$ associated with $E$ as
$h^{L}_{E}=\inf\Big\{h^{L}_{0}\text{ with psh potential }\mathrel{\mathop{\ordinarycolon}}h^{L}_{0}\geq h^{L}\text{ over }E\Big\}.$ (1.9)
When $L$ is ample, the non-pluripolarity of $E$ assures that the metric $h^{L}_{E}$ has a bounded potential, cf. [38, Theorem 9.17]. When $L$ is merely big, the potential of $h^{L}_{E}$ might have singularities. In either case, the lower semi-continuous regularization $h^{L}_{E*}$ of $h^{L}_{E}$ is a metric with a psh potential, and it has the minimal singularities, cf. [6, §1.2] for the necessary definitions.
Recall that the equilibrium measure associated with a non-pluripolar $(K,h^{L})$ is defined as
$\mu_{\mathrm{eq}}(K,h^{L})\mathrel{\mathop{\ordinarycolon}}=\frac{1}{{\rm{vol}}(L)}c_{1}(L,h^{L}_{K*})^{n}.$ (1.10)
Above, ${\rm vol}(L)$ denotes the volume of the line bundle $L$, defined by
${\rm{vol}}(L)\mathrel{\mathop{\ordinarycolon}}=\lim_{k\to\infty}\frac{n!}{k^{n}}\dim H^{0}(X,L^{\otimes k}),$ (1.11)
a limit that exists by Fujita’s theorem [35] and is strictly positive under our assumption that $L$ is big. The positive $(1,1)$-current $c_{1}(L,h^{L}_{K*})$ is defined as $c_{1}(L,h^{L}_{K*})\mathrel{\mathop{\ordinarycolon}}=c_{1}(L,h^{L}_{0})+\frac{\sqrt{-1}}{2\pi}\partial\bar{\partial}\phi_{K}$, where $h^{L}_{K*}=h^{L}_{0}\cdot\exp(-\phi_{K})$ and $h^{L}_{0}$ is an arbitrary smooth metric on $L$. It is generally non-smooth by the above discussion. When $L$ is ample, the top wedge power $c_{1}(L,h^{L}_{K*})^{n}$ can be defined in the sense of Bedford-Taylor [2], since $h^{L}_{K*}$ admits a bounded potential. However, as we assume $L$ is merely big, the wedge product $c_{1}(L,h^{L}_{K*})^{n}$ appearing in (1.10) is interpreted via the non-pluripolar product developed by Boucksom-Eyssidieux-Guedj-Zeriahi [17], which generalizes the Bedford-Taylor construction. In this case, $\mu_{\mathrm{eq}}(K,h^{L})$ is a probability measure on $X$ supported in $K$, see [6, Proposition 1.10]. The following result settles a question left open by Zelditch [64, after Theorem 2.5] and provides a refinement of Theorem 1.1, albeit under an additional assumption.
Theorem 1.3.
For any Bernstein-Markov measure $\mu$ with a non-pluripolar support, the measures $\mu_{k}^{\rm{Berg}}$ converge weakly, as $k\to\infty$, to the measure $\Delta_{*}\mu_{\mathrm{eq}}(K,h^{L})$, where $\Delta\mathrel{\mathop{\ordinarycolon}}X\to X\times X$ is the diagonal embedding.
Remark 1.4.
As we shall explain in Section 2.2, Theorem 1.3 refines the convergence of the diagonal Bergman measures, due to Berman-Boucksom-Witt Nyström [5] and Bloom-Levenberg [12], [13]. We, however, stress out that our proof depends on [5], and the only novel aspect of Theorem 1.3 is the proof of the asymptotic concentration of the mass of $\mu_{k}^{\mathrm{Berg}}$ along the diagonal, i.e. Theorem 1.1. We also underline that Theorem 1.1 can be established rather easily for measures satisfying the Bernstein-Markov assumption, see Section 2.2.
We state now a complementary form of the localization principle, phrased in terms of the diagonal Bergman measure, which is a probability measure on $X$ defined as
$\mu_{k}^{\rm{Berg},\Delta}\mathrel{\mathop{\ordinarycolon}}=\frac{1}{n_{k}}B_{k}(x,x)\cdot d\mu(x).$ (1.12)
The same heuristic discussed after (1.3) suggests that although the diagonal Bergman kernel is a global invariant of the measure, variations in the weight of the measure should affect the asymptotic value of the diagonal Bergman kernel at a given point only through the local change of the weight at that point, rather than globally. This observation is formalized in our next theorem, for which we fix two Borel measures $\mu_{1}$ and $\mu_{2}$ on $X$, so that for a certain continuous $f\mathrel{\mathop{\ordinarycolon}}X\to\real$, $\mu_{2}=\exp(f)\cdot\mu_{1}$. We denote by $\mu_{k,i}^{\rm{Berg},\Delta}$, $k\in\mathbb{N}$, $i=1,2$, the associated diagonal Bergman measures.
Theorem 1.5.
For $\mu_{1}$ and $\mu_{2}$ as above, the sequence of measures $\mu_{k,1}^{\rm{Berg},\Delta}-\mu_{k,2}^{\rm{Berg},\Delta}$ converges weakly to the zero measure, as $k\to\infty$, if $\mu_{1}$ or $\mu_{2}$ do not give full mass to pluripolar subsets.
Remark 1.6.
It is natural to ask whether Theorem 1.5 admits a proof that does not rely on Theorem 1.1. If so, Theorem 1.1 could in turn be derived by the methods of Section 2.2.
@@
Our primary motivation for the study of the localization phenomena of the Bergman kernel lies in the theory of Toeplitz operators. Let us recall the necessary notations: for $f\in\mathscr{C}^{0}(X,\real)$, $k\in\mathbb{N}^{*}$, we define the Toeplitz operator $T_{k}(f)\in{\text{End}{(H^{0}(X,L^{\otimes k}))}}$ as $T_{k}(f)\mathrel{\mathop{\ordinarycolon}}=B_{k}\circ M_{k}(f)$, where $B_{k}\mathrel{\mathop{\ordinarycolon}}\mathscr{C}^{0}(X,L^{\otimes k})\to H^{0}(X,L^{\otimes k})$ is the orthogonal (Bergman) projection to $H^{0}(X,L^{\otimes k})$ with respect to (1.1), and $M_{k}(f)\mathrel{\mathop{\ordinarycolon}}H^{0}(X,L^{\otimes k})\to\mathscr{C}^{0}(X,L^{\otimes k})$ is the multiplication map by $f$.
By considering the products $\langle T_{k}(f)s_{1},s_{2}\rangle_{{\textrm{Hilb}}_{k}(h^{L},\mu)}$ for arbitrary $s_{1},s_{2}\in H^{0}(X,L^{\otimes k})$, we see that $T_{k}(f)$ depends solely on the restriction of $f$ to $K$. As by Tietze-Urysohn-Brouwer extension theorem, any function from $\mathscr{C}^{0}(K,\real)$ admits an extension to $\mathscr{C}^{0}(X,\real)$, we can thus extend the definition of $T_{k}(f)$ for any $f\in\mathscr{C}^{0}(K,\real)$.
@@
One of the key features of Toeplitz operators is their asymptotic algebra property: for $f,g\in\mathscr{C}^{0}(K,\real)$, the composition $T_{k}(f)\circ T_{k}(g)$ is asymptotically close to $T_{k}(f\cdot g)$, as $k\to\infty$. This phenomenon appears to have been first observed by Grenander-Szegő in [36, §7.4] in the context of certain orthogonal polynomials, see Section 4.1. It was later established by Bordemann-Meinrenken-Schlichenmaier [15] when $\mu$ is a volume form, $L$ is ample and $h^{L}$ is smooth and positive; see also Ma-Marinescu [44], [45]. The following result shows that, somewhat unexpectedly, this phenomenon holds in great generality.
Theorem 1.7.
@@
For any Borel measure $\mu$ which does not give full mass to pluripolar subsets, the space of Toeplitz operators verifies the asymptotic algebra property. In other words, for any $f,g\in\mathscr{C}^{0}(K,\real)$, $p\in[1,+\infty[$, we have
@@
Apart from the case when $\mu$ is a volume form, $L$ is ample and $h^{L}$ is smooth and positive, the author is not aware of any general condition on $\mu$ for which the $p$-Schatten norm above could be replaced by the operator norm. Due to possible connections to symplectic geometry in light of the foundational works [15] and Ma-Marinescu [45], it would also be interesting to classify measures $\mu$, which allow for a refinement of (1.13) to include the lower-order terms in $k$, and ultimately to analyze the commutator $\frac{1}{k}[T_{k}(f),T_{k}(g)]$.
As an application of Theorem 1.7, we get the following equidistribution result for the spectral measures of Toeplitz operators.